Given α is non-income tax, β is income tax, δ is marginal propensity to consume, γ is autonomous consumption, T (tax), Y (national income), I0 (investment) and G0 (government expenditure).

a) Formulate the equations needed to find the reduced form of equilibrium income (Ye).
1
Expert's answer
2020-12-23T04:43:18-0500

The equilibrium level of income refers to when an economy or business has an equal amount of production and market demand. Most simply, the formula for the equilibrium level of income is when aggregate supply (AS) is equal to aggregate demand (AD), where AS = AD. Adding a little complexity, the formula becomes Y(e) = C + IO + GO, where Y(e) is aggregate income, C is consumption, IO is investment expenditure, and GO is government expenditure. Using this formula, an analyst can observe how a change in any of the factors will impact the level of income. For example, if government spending increases, and all other expenditures stay constant, the level of aggregate income must also increase to maintain the equilibrium level of income.

Y(E) = C + IO + GO

C = y + δ(Y - T) where, y>0.0<δ<1

T = α + βY where α>0.0<β<1